If the cost of producing each unit of product A is 50 Php and the selling price for each unit is 150 Php. By using the revenue and cost functions method, how much profit (loss) is earned by the business if it is sold 500 units of Product A?
Answer by ikleyn(52812)

.
If the cost of producing each unit of product A is 50 Php and the selling price for each unit is 150 Php.
By using the revenue and cost functions method, how much profit (loss) is earned by the business if it is sold 500 units of Product A?
~~~~~~~~~~~~~~~~~~

Cost producing function is  C(x) = 50x, where x is the number of items produced.


Revenue function is  R(x) = 150x,  where x is the number of items sold.


The profit is the difference  P(x) = R(x) - C(x).


In your problem, the profit is


    P(500) = 500*150 - 500*50 = 500*(150-50) = 500*100 = Php 50,000.    ANSWER

Solved.